OpenStudy (anonymous):

hi..can anyone solve this? please lim[x:+∞,ln(x)-(1)/(x^(2))]

OpenStudy (anonymous):

Is this your question? \[\large \lim_{x \rightarrow +\infty} \frac{\ln(x)-1}{x^2}\]

OpenStudy (anonymous):

yes :)..how did u do that?

OpenStudy (anonymous):

Apply l'hopital rule because it is infinity/infinity form

OpenStudy (anonymous):

Do you know l'hopital rule ??

OpenStudy (anonymous):

to tell the truth i dont know this..but i'll google it :) Thank you very much for the help ^_^

OpenStudy (anonymous):

It is very easy. Differentiate the numerator and differentiate the denomintor separately. That's all

OpenStudy (anonymous):

*denominator

OpenStudy (anonymous):

yeah..i got it ..thanks again :) i have a math test tomorrow and i have some excercises i do not understand.

OpenStudy (anonymous):

OK :)

OpenStudy (anonymous):

Do have a look here too. Very important before using l'hopital rule, read this http://www.math.tamu.edu/~fulling/coalweb/lhop.htm